# WildFireCan-MMD: A Multimodal dataset for Classification of User-generated Content During Wildfires in Canada

### v1

@@ Abstract
Rapid information access is vital during wildfires, yet traditional data sources are slow and costly. Social media offers real-time updates, but extracting relevant insights remains a challenge. We present WildFireCan-MMD, a new multimodal dataset of X posts from recent Canadian wildfires, annotated across 13 key themes. Evaluating both Vision Language Models and custom-trained classifiers, we show that while zero-shot prompting offers quick deployment, even simple trained models outperform them when labelled data is available, by up to 23%. Our findings highlight the enduring importance of tailored datasets and task-specific training. Importantly, such datasets should be localized, as disaster response requirements vary across regions and contexts.
@@ 1 Introduction
Large-scale analysis of online interactions, using machine learning algorithms has enabled critical applications such as disaster response Fauzi (2023); Kumar et al. (2024b), mental health monitoringGhanadian et al. (2024); Dalal et al. (2024), detecting and combating hate speech in digital environments Kiritchenko et al. (2021); Kumar et al. (2024a), and sentiment and behavioral trend analysis Ma et al. (2024); Curto et al. (2024). Among these applications, analysis of social media data during natural disasters has gained increasing importance in recent years. Such data contains crucial details such as reports of casualties, infrastructure damage, urgent needs of affected individuals, and information about missing persons, among other critical information Houston et al. (2015); Alam et al. (2018). Using the shared information effectively can facilitate humanitarian aid efforts for more efficient management and faster recovery. This valuable data, however, is mixed in with noise, such as humorous content, advertisements, and other irrelevant information. Therefore, there is a critical need to develop machine learning models trained for sifting through this information.
@@
To foster research in this area, Alam et al. (2018) have developed the CrisisMMD dataset to harness multimodal social media data for disaster response and humanitarian aid. This dataset was collected from Twitter and included earthquakes, hurricanes, wildfires, and floods occurring in 2017. However, the classes of CrisisMMD are generalized across all types of disasters, meaning they are not tailored to specific disaster scenarios. As a result, the classification scheme is broad, and certain nuances or unique information that may be relevant to a particular type of disaster could be overlooked. Various information, such as the nature of the damages and the type of assistance required, often differs depending on the specific disaster and its geographic context.
To address this limitation, we focus our work on gathering social media data specifically related to wildfires in Canada. Wildfires are becoming increasingly frequent and severe in this country due to the adverse effects of climate change. Although the CrisisMMD dataset does contain wildfire data from the California wildfires of 2017, the proportion of this type of data compared to the rest is relatively small, only comprising about %12 of the total dataset and specific to the US context. Therefore, wildfires in Canada remain underexplored in social media analysis research. Upon inspecting our wildfire-specific social media dataset, we observed that existing categorization schemes do not capture the specificity of wildfire-related social media content. To address this, we devised thirteen tailored categories reflecting the unique information needs and communication patterns of Canadian wildfire contexts.
Then, we explored the feasibility of training multimodal classification models for triaging tweets during wildfires. In recent years, the research community has extensively worked on developing multimodal machine learning models for processing and categorizing social media data during natural disasters such as wildfires, hurricanes, and floods Ofli et al. (2020); Dwarakanath et al. (2021); Algiriyage et al. (2022); Basit et al. (2023); Koshy and Elango (2023); Hossain et al. (2022); Rezk et al. (2023); Khattar and Quadri (2022). Almost all of these works use the CrisisMMD dataset for their research, highlighting the scarcity of datasets on this topic.
To Benchmark the available models on WildFireCan-MMD, we evaluate the zero-shot classification capabilities of VLMs. Specifically, we prompt LLaVA (open-source) and gpt-4o-mini (proprietary) for label generation, and benchmark their performance against custom-trained multimodal classifiers using concatenative early and late fusion of RoBERTa and ViT modelsOfli et al. (2020); Basit et al. (2023). While VLMs offer a fast, label-free alternative, our results show that fine-tuned fusion models significantly outperform them. The full set of labeled data (containing the tweet IDs and labels), as well as the experimental code is available at WildFireCanMMD-Github.  [^1] Please contact the authors for more information about data sharability.. The contributions of this study are summarized as follows:
  • We developed a 13-category taxonomy through multimodal topic modleing of wildfire-related user-generated content.
@@
  • We introduced WildFireCan-MMD, the first Canada-specific multimodal wildfire dataset, comprising 4,688 posts from X in 2022–2024, labeled based on our taxonomy.
  • Uisng WildFireCan-MMD as benchamrk, we evaluate zero-shot VLMS on this task, finding that while closed-source models outperform open ones, both fall short without fine-tuning, which highlights the need for curated training data. Fine-tuned classifiers outperform zero-shot VLMs significantly; our best model achieves an F-score of 0.83, exceeding gpt-4o-mini by 23%.
@@ 2 Background and Literature Review
In this section, we review the literature on social media analysis for disaster response beyond the machine learning field, as well as the literature on the technical aspects of this project.
@@ 2.1 Social media analysis for disaster response
The growing interest among researchers in the role of social media during disasters has led to significant contributions in this field. Reference Fauzi (2023) conducted a bibliometric analysis to explore the use of social media in disaster management. This work provides a detailed review of the field, identifying key publications and future research directions, such as the implications of climate emergencies, integrating big data analytics, and managing knowledge in disaster contexts. Similarly, Muniz-Rodriguez et al. (2020) focused on the public health aspects of social media use during emergencies, noting its effectiveness in spreading emergency warnings and assessing needs after a disaster. Furthermore, Algiriyage et al. (2022) offered a systematic review of the application of deep learning in disaster response. They discuss the successes and challenges faced in this area, propose guidelines for future research, and consider the importance of using multimodal data in disaster scenarios. These contributions highlight the significant and varied ways social media can support disaster response efforts. Also, Acikara et al. (2023) explored how social media analytics can improve disaster response by enabling the rapid collection of time-sensitive information from affected areas. They noted the value of social media users as "citizen sensors" for quickly gathering essential data. Their review suggests that using social media analytics in disaster management, particularly for natural and human-made hazards, is a developing area that emphasizes the importance of human-centric insights, such as collective intelligence and location awareness. They also identify a gap in research on managing transport networks in real-time after disasters, which could enhance community resilience and sustainability. In other works, (Dwarakanath et al. (2021), Bukar et al. (2020), Mundottukandi et al. (2024)) discussed the crucial role of social media in managing crises, pointing out how platforms like Facebook and Twitter support communication and coordination among affected individuals and communities. Their reviews offer guidance for selecting research topics and aiding emergency teams in making informed decisions.
Our work is motivated by this literature and focuses on the use of machine learning models to triage social media data to potentially be used in all phases of disaster management.
@@ 2.2 Multimodal classification for disaster response
Expanding beyond text-based analysis of social media content, recent studies have emphasized multimodal analysis. We review this literature based on the available datasets and previously explored algorithms.
Datasets: Reference Dong et al. (2021) aimed to improve disaster relief by analyzing social media data to understand public attitudes and demand for relief supplies across different disaster types, using a textual Twitter dataset. They apply both qualitative and quantitative methods, including eight machine learning models, to assess public perception, offering insights for more effective disaster management. Also, Ochoa and Comes (2021) introduced a machine learning model that uses multimodal data, including satellite imagery and social media, to predict housing and shelter needs following disasters, demonstrating the model’s accuracy in forecasting such requirements. The lack of labeled social media imagery data for disaster response has been addressed by Alam et al. (2018), who created CrisisMMD, a comprehensive multimodal dataset from Twitter covering various natural disasters. This dataset, with thousands of manually annotated tweets and images from events like earthquakes, hurricanes, wildfires, and floods, provides valuable annotations for crisis management tasks. The dataset has been extensively utilized in research, shedding light on the complex communication patterns during disasters and the importance of integrating textual and visual data to better understand social media’s role in disaster contexts. The work by Ofli et al. (2020) split the problem into two distinct tasks. ’Task 1’ being informative or non-informative, to determine at a high level whether a given tweet is useful for humanitarian aid purposes. ’Task 2’ being the humanitarian categories or classes which the tweets are sorted into, for which they had a total of eight.
@@
While previous works used a generalized scheme of annotations across various disasters, our work develops a wildfire-specific taxonomy of thirteen classes and introduces, WildFireCan-MMD, a Canadian dataset annotated with these labels.
@@
Algorithms: Previous work by Ofli et al. (2020) introduced a joint representation learning from two parallel deep learning architectures where one architecture represents the text modality and the other represents the image modality. Image and text modalities are vectorized through VGG16 and Word2Vec+CNN architectures, respectively, followed by a dense layer for prediction. The work by Basit et al. (2023), expands on Ofli et al. (2020) by experimenting with multiple recent transformer structures, for both the text and image modalities, and compared them against the baseline reported in the work by Ofli et al. (2020). They found that using a combination of AlBERTLan et al. (2020) as a text embedder, and DeITTouvron et al. (2021) as an image embedder, provided the best results for the eight-class classification task. Reference Koshy and Elango (2023) implemented further experiments with text and image transformers and showed improvements by using bi-LSTMs instead of the traditional LSTM. Furthermore, they have noted that multiplicative fusion strategies produce better results on a classification task when combined with the previously described methods. Specifically, they demonstrated that using the fine-tuned RoBERTaLiu (2019) model for text, Vision Transformer model for image, biLSTM+attention mechanism with multiplicative fusion, for prediction, achieves the best results for the eight-class classification task described by Alam et al. (2018).
We build on the previous work and explore the existing structures on the newly annotated data, WildFireCan-MMDintroduced in this work.
2.3 Large Vision-Language Models
Recent advances in multimodal LLMs have enabled models to process image and text jointly, expanding their utility across a range of complex tasks. A comprehensive survey by Zhang et al. (2024) outlines the design, training strategies, and benchmark performance of 26 MM-LLMs, while Li et al. (2022) categorizes the evolution of VLMs into task-specific models, vision-language pretraining approaches, and those leveraging large-scale, weakly labeled image-text data. These VLMs now support zero- and few-shot generalization across diverse domains.
In this work, we evaluate gpt-4o-mini and LLaVA as zero-shot classifiers on the WildFireCan-MMD dataset. gpt-4o-mini, OpenAI’s most recent transformer-based VLM at the time Achiam et al. (2023), is closed-source and accessible only via API, limiting transparency around its architecture and training. In contrast, LLaVA Liu et al. (2023) is fully open-source and trained using GPT-4-generated responses paired with user-generated prompts and CLIP-based visual features Radford et al. (2021). It employs visual instruction tuning to align vision and language modalities and has shown strong performance on multimodal instruction-following benchmarks. The release of models like LLaVA reflects broader efforts to build high-performing, publicly accessible VLMs in response to the closed nature of commercial systems like GPT-4.
@@ 3.1 Data Collection
We used X’s Pro-Tier paid API to gather a diverse dataset of posts containing information, updates, and discussions related to the British Columbia/Alberta wildfires of 2022 and 2023, and the Jasper wildfires of 2024. This dataset contains tweets between the months of April and June for each year and consists of 4,688 unique image-text pairs. We collected this data using the Tweepy library and conducted a hashtag-based search, using hashtags we found to yield good results through experimentation. This was a crucial part of our data collection as we wanted to ensure we gathered tweets originating from Canada, that were also on topic. Due to not many users in Canada having their specific geo-location turned on for Twitter, we couldn’t simply scrape by region, or very little data would be returned. As a result, we had to adopt a hashtag-based search. Due to the uniqueness of the hashtags, which are only really used in Canada during wildfires, this yielded a dataset of relevant tweets with mostly users from the specified region. The dataset includes the timestamp of when the tweet was posted, the user-provided location, and, in some cases, the specific region or county from which the tweet originated. Table A.1 presents the full queries used for data collection.
3.2 Developing a Taxonomy
@@
To examine the content of the tweets and discover the hidden, latent semantics of our data, we used the BERTopic to perform multimodal unsupervised topic modeling Grootendorst (2022); Falkenberg et al. (2022); Egger and Yu (2022); Medvecki et al. (2023). We then manually inspected the topics to identify the categories. In doing so, we considered the stakeholders who will be using this data. Implementation details of the BERTopic clustering are provided in Appendix B. In the end, we selected 13 classes which were important and recurrent enough to be considered categories for this dataset. The full list of categories and their descriptions is given in Table 1.
@@
The selection of classes was guided by two primary considerations: the frequency of occurrence observed through manual annotation and the perceived relevance of classes to stakeholders in disaster scenarios. For example, ‘Warnings & Status Updates’, ‘Reports of Actions of Responders’, and ‘Infrastructure’ could be useful to first responders and emergency personnel. The ‘Evacuees,’General Information,’ and ‘Preparedness’ categories could be useful to the general public to be informed about the events happening around their locality. The ‘Political’ category could be useful to authorities to gauge public opinion on events surrounding the disaster.
@@
Table 1: Categories of WildFireCan-MMD, their descriptions and the size of each class.
Category Description Class Size
@@ 3.3 Data Labeling
We labeled the collected dataset according to the taxonomy of categories presented in Table 1. Labelling was done in-house by one of the authors, using Label Studio [^2] https://labelstud.io/. Since this task is not subjective, we did not consider multi-annotator labeling. Label Studio was chosen for its ability to search through the text fields of the dataset for specific keywords, rather than running through the dataset in sequence, which made labeling easier. For example, this allowed us to search for ’air quality’ and see all posts containing these keywords. This allowed us to focus on one category at a time and get the majority of a category labeled before moving on to another. When searching by keywords was no longer possible, labeling in sequence resumed.
@@
Examples of labeled tweets are presented in Figure 1. Figure 2 displays the class distribution in WildFireCan-MMD. Also, for training and testing the custom classifiers, given in Section 4, we split the data into an 80/20 train/test split using stratified sampling, to ensure similar distributions of classes in train and test subsets. The size of labeled classes are presented in Table 1.
Figure 2: Class distribution of WildFireCan-MMD.

### v2

@@ Abstract
Rapid information access is vital during wildfires, yet traditional data sources are slow and costly. Social media offers real-time updates, but extracting relevant insights remains a challenge. We present WildFireCan-MMD, a new multimodal dataset of X posts from recent Canadian wildfires, annotated across twelve key themes. Evaluating both vision-language models and custom-trained classifiers, we show that while zero-shot prompting offers quick deployment, even simple trained models outperform them when labelled data is available. Our best-performing transformer-based fine-tuned model reaches 83% f-score, outperforming gpt4 by 23%. Also, we demonstrate how this model can be used to uncover trends during wildfires. Our findings highlight the enduring importance of tailored datasets and task-specific training. Importantly, such datasets should be localized, as disaster response requirements vary across regions and contexts.
@@ 1 Introduction
Large-scale analysis of online interactions, using machine learning algorithms has enabled critical applications such as disaster response Fauzi (2023); Kumar et al. (2024b), mental health monitoring Ghanadian et al. (2024); Dalal et al. (2024), detecting and combating hate speech in digital environments Kiritchenko et al. (2021); Kumar et al. (2024a), pandemic discourse analysis Imran et al. (2022) and sentiment and behavioral trend analysis Ma et al. (2024); Curto et al. (2024). Among these applications, analysis of social media data during natural disasters has gained increasing importance in recent years Alam et al. (2018); Ofli et al. (2020); Alam et al. (2021); Dwarakanath et al. (2021); Algiriyage et al. (2022); Hossain et al. (2022); Khattar and Quadri (2022); Basit et al. (2023); Koshy and Elango (2023); Rezk et al. (2023). Such data contains crucial details such as reports of casualties, infrastructure damage, and urgent needs of affected individuals, among other critical information. The effective use of the shared information can facilitate humanitarian aid efforts for more efficient management and faster recovery Houston et al. (2015). This valuable data, however, is mixed in with noise, such as humorous content, advertisements, and other irrelevant information. Therefore, there is a critical need to develop an automated method for sifting through this information.
@@
To foster research in this area, Alam et al. (2018, 2021) developed two multimodal datasets, CrisisMMD and HumAID, both collected from Twitter, including several disasters such as earthquakes and hurricanes. However, the classes of these datasets are generalized across all types of disasters, meaning they are not tailored to specific disaster scenarios. As a result, the classification scheme is broad, and certain nuances or unique information that may be relevant to a particular type of disaster could be overlooked. Also, while the CrisisMMD dataset does contain wildfire data from the California wildfires of 2017, the proportion of this type of data compared to the rest is relatively small, only comprising about %12 of the total dataset and specific to the US context. The HumAID dataset by Alam et al. (2021) also contains wildfire data from Canada, but only a couple of thousand samples from 2016.
Therefore, wildfires in Canada remain underexplored in social media analysis research. To address this limitation, we focus our work on gathering social media data specifically related to Canadian wildfires, as they are becoming increasingly frequent and severe due to the adverse effects of climate change.
Upon initial inspection of our Canadian wildfire-specific social media dataset, we observed that existing categorization schemes do not capture the specificity of this data. To address this, we devised a novel taxonomy of twelve categories reflecting the unique information needs and communication patterns of Canadian wildfire contexts. Using our developed taxonomy, we manually labeled 4,688 wildfire-related tweets from Canada and created WildFireCan-MMD, a new multimodal dataset of X posts from recent Canadian wildfires.
To benchmark the available classification models on WildFireCan-MMD, we evaluated the zero-shot classification capabilities of several VLMs, as well as the feasibility of training custom classifiers. Our best-performing classifier is a fine-tuned model, which uses RoBERTa Liu (2019) and Vision Transformer (ViT) models Ofli et al. (2020); Basit et al. (2023) to jointly process text and image parts of tweets. While WildFireCan-MMD is sourced from data posted in 2022-2024, we collected an additional set of unlabeled tweets across seven years (2018-2025). We then used our best-performing classifier to analyze this data and uncover trends and patterns of social media use in Canada. The full set of labeled data (containing the tweet IDs and labels), as well as the experimental code, is available at WildFireCanMMD-Github [^1] Please contact the authors for more information about data sharability., under CC-BY-NC license. The contributions of this study are summarized as follows:
@@
  • We devised a twelve-category taxonomy through multimodal topic modeling of Canadian wildfire-related user-generated content, and introduced WildFireCan-MMD, the first Canada-specific multimodal wildfire dataset, comprising 4,688 posts from X in 2022–2024, manually labeled based on our taxonomy.
  • We compared custom multimodal transformer models trained on WildFireCan-MMD with several VLMs and classical methods. Fine-tuned classifiers outperform zero-shot VLMs significantly; our best model achieves an F-score of 0.83, exceeding gpt-4o-mini by 23%.
  • We collected an additional 46,279 posts from X, spanning across seven years, used our trained classifier to label them, and reported the uncovered trends and patterns of social media use during wildfires in Canada.
@@ 2.1 Social media analysis for disaster response
Significant research has been conducted on the role of social media during disasters. Fauzi (2023) conducted a bibliometric analysis of social media use in disaster response and offered a comprehensive overview of key studies and emerging research directions. Acikara et al. (2023) highlighted the potential of social media analytics in disaster response by leveraging users as “citizen sensors” to rapidly gather critical, time-sensitive information. Muniz-Rodriguez et al. (2020) focused on the public health aspects of social media use during emergencies, noting its effectiveness in spreading emergency warnings and assessing needs after a disaster. In other works, Dwarakanath et al. (2021), Bukar et al. (2020), and Mundottukandi et al. (2024) discussed the crucial role of social media in managing crises, pointing out how platforms like Facebook and Twitter support communication and coordination among affected individuals and communities. Furthermore, in a systematic review of the application of deep learning in disaster response, Algiriyage et al. (2022) discussed the successes and challenges faced in this area, proposed guidelines for future research, and highlighted the importance of using multimodal data in disaster scenarios.
Our work is motivated by this literature and focuses on using machine learning models to triage Canada-specific multimodal social media data to inform disaster response and management.
@@ 2.2 Multimodal classification for disaster response
We review this literature based on the available datasets and previously explored algorithms.
Datasets: While many works utilize textual datasets sourced from social media to classify posts during disasters Olteanu et al. (2014); Dong et al. (2021); Ochoa and Comes (2021), it is now generally considered that multimodal approaches offer stronger results Ofli et al. (2020); Abavisani et al. (2020); Ramirez et al. (2025).
Alam et al. (2018) addressed the shortage of multimodal labeled social media data for disaster response by introducing CrisisMMD, a widely used dataset containing annotated tweets and images from various natural disasters. Later extended by datasets like HumAID Alam et al. (2021), CrisisMMD has become a foundational resource for studying disaster-related communication patterns and the combined use of text and images. It remains the most prominent dataset in this domain, used in many recent studies Basit et al. (2023); Koshy and Elango (2023); Shetty et al. (2024).
@@
While previous works such as CrisisMMD used a generalized scheme of annotations across various types of different disasters, our work develops a wildfire-specific taxonomy of twelve categories and introduces WildFireCan-MMD, a new Canadian wildfire dataset.
@@
Algorithms: Previous work by Ofli et al. (2020) introduced a joint representation learning from two parallel deep learning architectures where one architecture represents the text modality and the other represents the image modality. Image and text modalities are vectorized through VGG16 and Word2Vec+CNN architectures, respectively, followed by a dense layer for prediction. The work by Basit et al. (2023), expands on Ofli et al. (2020) by experimenting with multiple recent transformer structures, for both the text and image modalities, and compared them against the baseline reported in the previous work. They found that using a combination of AlBERT Lan et al. (2020) as a text embedder, and DeIT Touvron et al. (2021) as an image embedder, provided the best results for the eight-class classification task. Koshy and Elango (2023) implemented further experiments with text and image transformers and showed improvements by using bi-LSTMs instead of LSTMs. Furthermore, they have noted that multiplicative fusion strategies produce better results on a classification task when combined with the previously described methods. Specifically, they demonstrated that using the fine-tuned RoBERTa Liu (2019) model for text, Vision Transformer model for image, biLSTM+attention mechanism with multiplicative fusion, for prediction, achieves the best results for the eight-class classification task described by Alam et al. (2018). We build on the previous work and explore the existing structures on the newly annotated data, WildFireCan-MMD.
In addition to fine-tuned structures, we explore using VLMs in a zero-shot setting Zhang et al. (2024); Li et al. (2022). Specifically, we evaluate three VLMs, GPT-4 Achiam et al. (2023), LLaVA Liu et al. (2023), and SmolVLM Marafioti et al. (2025), to assess performance across a range of capabilities and computational requirements. GPT-4 serves as a state-of-the-art, closed-source benchmark, while LLaVA and SmolVLM are open-source models that use visual instruction tuning and are optimized for varying levels of hardware, reflecting ongoing efforts to build accessible, high-performing multimodal systems.
@@ 3.1 Data Collection
To construct our dataset, we utilized X’s Pro-Tier paid API to collect a diverse set of posts surrounding the British Columbia and Alberta wildfires of 2022 and 2023, as well as the Jasper wildfires of 2024. The dataset comprises 4,688 unique image-text pairs gathered from tweets posted between April and June of each respective year. Data collection was conducted using a hashtag-based search strategy. Table A.1 presents the full queries used for data collection. To curate the hashtags, we started with seed hashtags and used an iterative process of adding new hashtags by collecting sample tweets. This approach was necessary due to the limited availability of geotagged content, as most users on the platform do not enable precise geo-location. Instead, we identified and utilized hashtags that, through iterative experimentation, were found to be effective indicators of relevant content originating from Canada during wildfire events. The specificity of these hashtags ensured that the collected tweets were not only on-topic but also predominantly from users in the targeted geographic regions. The dataset includes the timestamp of when the tweet was posted, the user-provided location, and, in some cases, the specific region or county from which the tweet originated.
3.2 Taxonomy Development
@@
To examine the content of the tweets and discover the hidden latent semantics of our data, we used BERTopic to perform multimodal unsupervised topic modeling Grootendorst (2022); Falkenberg et al. (2022); Egger and Yu (2022); Medvecki et al. (2023). Further explanation of the BERTopic clustering method are provided in Appendix B.
Employing BERTopic on our dataset, we first let the algorithm determine the number of latent topics instead of setting number of topics to sort the data into. When the number of clusters were unspecified, BERTopic had created over 100 fine-grained topics. This was relatively large to manage, and difficult to determine a humanitarian category for each fine-grained topic. We gradually decreased the number of topics, and for each number, we manually inspected the representative visual samples of topics and keywords generated by BERTopic. Once we had reduced our topics to around 15, it became clearer which topics were the strongest in our dataset, and we were able to begin to see the humanitarian utility of some of these topics. Some examples of useful topics were posts about infrastructure, weather reports, and reports of actions of first responders. We also noticed some less useful posts, such as advertisements and political debates/complaints. Some examples of these topics, including representative images and corresponding keywords, generated by BERTopic, can be seen in Figure B.1. We then manually inspected the topics and selected 12 classes that were informative and consistently appeared across a varying maximum number of clusters to be considered designated categories for this dataset.
@@
The full list of categories and their descriptions is given in Table 1. We also include an “Other” to accommodate the examples that do not fit into any of these 12 categories. The selection of classes was guided by two primary considerations: the frequency of occurrence observed through manual annotation and the perceived relevance of classes to stakeholders in disaster scenarios. For example, ‘Warnings & Status Updates’, ‘Reports of Actions of Responders’, and ‘Infrastructure’ could be useful to first responders and emergency personnel. The ‘Evacuees,’General Information,’ and ‘Preparedness’ categories could be useful to the general public to be informed about the events happening around their locality. The ‘Political’ category could be useful to the media to gauge public opinion and policies on events surrounding the disaster.
@@
Table 1: Categories of WildFireCan-MMD, their descriptions and total count of each class.
Category Description Count
@@ 3.3 Data Labeling
We labeled the collected dataset according to the taxonomy of categories presented in Table 1. Over the course of labeling, we ensured that the data did not contain personal information. Labeling was carried out in-house by one of the authors using Label Studio [^2] https://labelstud.io/. To ensure the reliability of the labeling process, we first discussed the definitions and boundaries of each category among the research team and reached a shared understanding. Through this discussion, we determined that the task was not largely subjective, as the categories are well-defined and grounded in observable cues (e.g., specific topics, hazards, or references in the text or image). Therefore, we did not consider multi-annotator labeling necessary. Examples of labeled tweets are presented in Figure 1. The size of labeled classes are presented in Table 1.